# supplement

### v2

@@ 5.2.2 Several power iterations with truncation
(34) $v^{(p)}=A^{p}v^{(0)}+\sum_{i=1}^{p}A^{p-i}e^{(i)},$
where $e^{(i)}$ denotes a perturbation due to rank truncation at iteration $i$. In (34) each error vector $e^{(i)}$ is multiplied by $A^{p-i}$ so that as $p$ increases, truncation error from early iterations are amplified by large powers of $A$. Consequently, these errors become more aligned with $\psi_{1}$, mitigating their impact on convergence. In contrast, low-rank Krylov methods are more sensitive to large truncation errors in the early iterations. Such errors can cause significant deviations from the true Krylov subspace, often resulting in stagnation and poor convergence. This distinction highlights a key advantage of rank-truncated subspace iteration over rank truncated Krylov methods.
@@ 5.3 Subspace iteration
(35) $\tan\Theta(\mathcal{R}(V),\mathcal{R}(\Psi_{\leq m}))=\sigma_{1}(T_{V}),$
@@
(36) $T_{V}=\Psi_{>m}^{\text{T}}V\left(\Psi_{\leq m}^{\text{T}}V\right)^{-1},$
@@
(37) $\tan\Theta(\mathcal{R}(AV),\mathcal{R}(\Psi_{\leq m}))\leq\left|\frac{\lambda_{m+1}}{\lambda_{m}}\right|\tan\Theta(\mathcal{R}(V),\mathcal{R}(\Psi_{\leq m})).$
The proof of this theorem is provided in Appendix B.
@@ 5.3.1 Subspace iteration with truncation
(38) $V^{\prime}=Z\Phi+E^{\texttt{s}}=AV\Phi+E^{\texttt{mv}}\Phi+E^{\texttt{s}}\\ =AV\Phi+E,$
@@
(39) $\frac{\|E\|_{2}}{|\lambda_{m}|}<\left(\frac{1V}{\kappa_{2}(\Psi_{\leq m}^{\text{T}}V\Phi)}-\left|\frac{\lambda_{m+1}}{\lambda_{m}}\right|\right)\frac{\left\lVert\Psi_{\leq m}^{\text{T}}V\Phi\right\rVert_{2}\cdot\left\lVert\Psi_{>m}^{\text{T}}V\Phi\right\rVert_{2}}{\left\lVert V\Phi\right\rVert_{2}},$
@@
(40) $\Theta(\mathcal{R}(V^{\prime}),\mathcal{R}(\Psi_{\leq m}))<\Theta(\mathcal{R}(V),\mathcal{R}(\Psi_{\leq m}))$
@@
Proof 5.6.
The left-hand side of the inequality (40) is the largest singular value of the matrix $T_{V^{\prime}}$ which we write as
(41) $T_{V^{\prime}}$ $=\Psi_{>m}^{\text{T}}(AV\Phi+E)\left[\Psi_{\leq m}^{\text{T}}\left(AV\Phi+E\right)\right]^{-1}$
@@
(42) $\sigma_{1}(T_{V^{\prime}})$ $\leq\sigma_{1}(\Lambda_{>m}V_{>m}\Phi+E_{>m})\cdot\sigma_{1}\left((\Lambda_{\leq m}V_{\leq m}\Phi+E_{\leq m})^{-1}\right)$
@@
(43) $\sigma_{1}\left(\Lambda_{>m}V_{>m}\Phi+E_{>m}\right)$ $\leq\sigma_{1}\left(\Lambda_{>m}V_{>m}\Phi\right)+\left\lVert E_{>m}\right\rVert_{2}$
@@
(44) $\sigma_{m}\left(\Lambda_{\leq m}V_{\leq m}\Phi+E_{\leq m}\right)$ $\geq\sigma_{m}\left(\Lambda_{\leq m}V_{\leq m}\Phi\right)-\left\lVert E_{\leq m}\right\rVert_{2}$
@@
(45) $\sigma_{1}(T_{V^{\prime}})\leq\frac{|\lambda_{m+1}|\cdot\left\lVert V_{>m}\Phi\right\rVert_{2}+\left\lVert E_{>m}\right\rVert_{2}}{|\lambda_{m}|\cdot\sigma_{m}(V_{\leq m}\Phi)-\left\lVert E_{\leq m}\right\rVert_{2}}.$
@@
(46) $\sigma_{1}(T_{V})$ $=\sigma_{1}\left((V_{>m}\Phi)(V_{\leq m}\Phi)^{-1}\right)$
@@
(47) $\frac{|\lambda_{m+1}|\cdot\left\lVert V_{>m}\Phi\right\rVert_{2}+\left\lVert E_{>m}\right\rVert_{2}}{|\lambda_{m}|\cdot\sigma_{m}(V_{\leq m}\Phi)-\left\lVert E_{\leq m}\right\rVert_{2}}<\frac{\left\lVert V_{>m}\Phi\right\rVert_{2}}{\left\lVert V_{\leq m}\Phi\right\rVert_{2}}.$
@@
(48) $\left\lVert V_{\leq m}\Phi\right\rVert_{2}\left(|\lambda_{m+1}|\left\lVert V_{>m}\Phi\right\rVert_{2}+\left\lVert E_{>m}\right\rVert_{2}\right)$
$\qquad<\left\lVert V_{>m}\Phi\right\rVert_{2}\left(|\lambda_{m}|\sigma_{m}(V_{\leq m}\Phi)-\left\lVert E_{\leq m}\right\rVert_{2}\right),$
@@
(49) $\left\lVert V_{\leq m}\Phi\right\rVert_{2}\left\lVert E_{>m}\right\rVert_{2}+\left\lVert V_{>m}\Phi\right\rVert_{2}\ \left\lVert E_{\leq m}\right\rVert_{2}$
@@
(50) $\left\lVert V_{\leq m}\Phi\right\rVert_{2}\left\lVert E_{>m}\right\rVert_{2}+\left\lVert V_{>m}\Phi\right\rVert_{2}\ \left\lVert E_{\leq m}\right\rVert_{2}$ $\leq\left\lVert\begin{bmatrix}\left\lVert V_{\leq m}\Phi\right\rVert_{2}\\ \left\lVert V_{>m}\Phi\right\rVert_{2}\end{bmatrix}\right\rVert_{2}\left\lVert\begin{bmatrix}\left\lVert E_{>m}\right\rVert_{2}\\ \left\lVert E_{\leq m}\right\rVert_{2}\end{bmatrix}\right\rVert_{2}$
@@
(51) $\left\lVert E\right\rVert_{2}<\left(\frac{|\lambda_{m}|}{\kappa_{2}(V_{\leq m}\Phi)}-|\lambda_{m+1}|\right)\frac{\left\lVert V_{\leq m}\Phi\right\rVert_{2}\left\lVert V_{>m}\Phi\right\rVert_{2}}{\left\lVert V\Phi\right\rVert_{2}}$
where $\kappa_{2}(V_{\leq m}\Phi)=\sigma_{1}(V_{\leq m}\Phi)/\sigma_{m}(V_{\leq m}\Phi)$. Finally dividing the preceding inequality by $|\lambda_{m}|$ we arrive at (39).
@@
If the eigenvectors of interest $\psi_{1},\ldots,\psi_{m}$ have rank less than or equal to $\bm{r}$, the truncation error $E^{\texttt{mv}}$ resulting from the matrix-vector products in line 3 of the truncated subspace iteration (Algorithm 3) naturally decreases as the method converges. Similarly, the rank truncation in line 4 of the low-rank Rayleigh-Ritz procedure (Algorithm 4) introduces error that also decreases as the subspace iteration converges. This is because $v_{j}$’s converge to $\psi_{j}$’s, for $j=1,2,...,m$, which are naturally orthogonal for symmetric $A$. As a result, the linear combinations in the Rayleigh-Ritz step involve fewer terms of significant magnitude, thereby reducing truncation error.
6 Computational cost and implementation issues
In this section, we estimate the computational cost and discuss several practical issues that are important for efficient implementations of inexact subspace iteration.
@@ 6.2 Computing linear combination of low-rank TTs
In the Rayleigh-Ritz step (Algorithm 4), approximating each linear combination in line 4 using the TT-SVD approach involves applying a truncation after each addition, resulting in $m-1$ truncations. Each truncation has a computational cost of $\mathcal{O}(dnr^{3})$ and the ordering of the summands can affect the final result. Alternatively, approximating the entire sum directly on the tangent space of the TT manifold at $z_{k}$ requires a single orthogonal projection, which costs $\mathcal{O}(dnr^{3})$. For a large number of summands $m$, the tangent space projection becomes significantly more efficient.
@@ 7 Numerical results
In this section, we present several numerical examples to demonstrate the efficiency and effectiveness of the rank truncated subspace iteration and compare it with both the rank truncated Lanczos algorithm and the DMRG algorithm for computing several algebraically smallest eigenvalues. To initialize the linear transformation defining the Chebyshev polynomial (18) we estimate the largest eigenvalue using a few rank-truncated Lanczos iterations with safeguards as outlined in [39] and set $b$ equal to such approximation. We initialize $a=-1$ and update $a$ to be the largest Ritz value after each subspace iteration.
@@ 7.1 Test problems
    (52) $H=\sum_{j=1}^{L}-J\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+\sigma_{j}^{z}\sigma_{j+1}^{z}\right)-h\sigma_{j}^{z},$
@@
    In both cases, the Hamiltonian in (52) is formulated in the canonical polyadic (CP) decomposition format (2) and can be efficiently represented as a low-rank TT-matrix, also known as a matrix product operator (MPO).
@@
    (53) $\Delta=-\left(D\otimes I\otimes\cdots\otimes I+\cdots+I\otimes\cdots\otimes I\otimes D\right),$
@@ A Hamiltonian with a Hénon-Heiles potential can be written as
    (54) $H=-\frac{1}{2}\Delta+\frac{1}{2}\sum_{k=1}^{d}q_{k}^{2}+\mu\sum_{k=1}^{d-1}\left(q_{k}^{2}q_{k+1}-\frac{1}{3}q_{k+1}^{3}\right),$
@@ 7.2 Comparing subspace iteration and Lanczos method
We compared the rank truncated subspace iteration (Algorithm 2) and Lanczos method (Algorithm 1) for computing the first several smallest eigenvalues of each test problem (52)-(54). For (52) we set $L=10$ resulting in $H$ with size $2^{10}\times 2^{10}$, small enough to compute eigenvalues and eigenvectors using standard methods, which serve as benchmarks for evaluating the low-rank methods. Fig. 2 shows the error versus iteration for the first $5$ Ritz values computed by the low-rank subspace iteration and the low-rank Lanczos method. While the convergence Lanczos method stagnates around $10^{-3}$, the subspace iteration achieves convergence to machine precision. We also ran the Lanczos method without rank truncation to obtain the orthogonal Krylov basis vectors that the rank-truncated Lanczos method aims to approximate. In the right panel of Fig. 2, we plot the maximum of the TT-rank vector required to represent the first $20$ Krylov basis vectors with accuracy $10^{-10}$. After just two iterations, this rank exceeds $30$ indicating that these Krylov basis vectors can not be efficiently represented with small TT rank. Furthermore, Krylov methods are known to be sensitive to large errors in the early iterations, indicating that the stagnation in convergence arises from the inability to accurately represent the orthogonal Krylov basis vectors as low-rank TTs. In contrast, the subspace iteration does not encounter convergence issues due to rank truncation. Unlike the Krylov basis vectors, the ground state and first excited states can be effectively represented as low-rank TTs, as shown in the right panel of Fig. 2. Moreover the subspace iteration is robust to large truncation errors in the initial iterations.
@@ 7.3 Comparing subspace iteration and DMRG
We now compare rank-truncated subspace iteration with DMRG on spin-1/2 and spin-1 Heisenberg Hamiltonian (52) with $L=100$, no external magnetic field ($h=0$), $J=-1$ and periodic boundary conditions. We computed the ground state and first excited state energies (two algebraically smallest eigenvalues) using subspace iteration and DMRG method and plot the energies versus CPU-time in Fig. 3. In both cases we used a maximum rank of 100 for all vectors and both methods are initialized with the same random initial vectors. For the spin-1/2 system, the DMRG method and subspace iteration converge to the same eigenvalues with DMRG converging significantly faster. For the spin-1 system, DMRG also converges faster but gets stuck in a local minimum, a behavior that has been previously reported [37]. In contrast, subspace iteration achieves lower energies, with the ground state energy converging to the reference value.

### v3

@@ 4.1.2 Several power iterations with truncation
$v^{(p)}=A^{p}v^{(0)}+\sum_{i=1}^{p}A^{p-i}e^{(i)},$ (28)
where $e^{(i)}$ denotes a perturbation due to rank truncation at iteration $i$. In (28) each error vector $e^{(i)}$ is multiplied by $A^{p-i}$ so that as $p$ increases, truncation error from early iterations are amplified by large powers of $A$. Consequently, these errors become more aligned with $\psi_{1}$, mitigating their impact on convergence. In contrast, low-rank Krylov methods (see appendix A) are more sensitive to large truncation errors in the early iterations. Such errors can cause significant deviations from the true Krylov subspace, often resulting in stagnation and poor convergence. This distinction highlights a key advantage of rank-truncated subspace iteration over rank truncated Krylov methods.
@@ 4.2 Subspace iteration
$\tan\Theta(\mathcal{R}(V),\mathcal{R}(\Psi_{\leq m}))=\sigma_{1}(T_{V}),$ (29)
@@
$T_{V}=\Psi_{>m}^{\text{T}}V\left(\Psi_{\leq m}^{\text{T}}V\right)^{-1},$ (30)
@@
$\tan\Theta(\mathcal{R}(AV),\mathcal{R}(\Psi_{\leq m}))\leq\left|\frac{\lambda_{m+1}}{\lambda_{m}}\right|\tan\Theta(\mathcal{R}(V),\mathcal{R}(\Psi_{\leq m})).$ (31)
The above theorem is a well-known result, see, e.g., [29], we also provide a proof in appendix D that is consistent with the notation in the present paper.
@@ 4.2.1 Subspace iteration with truncation
$V^{\prime}=Z\Phi+E^{\texttt{s}}=AV\Phi+E^{\texttt{mv}}\Phi+E^{\texttt{s}}\\ =AV\Phi+E,$ (32)
@@
$\frac{\|E\|_{2}}{|\lambda_{m}|}<\left(\frac{1}{\kappa_{2}(\Psi_{\leq m}^{\text{T}}V\Phi)}-\left|\frac{\lambda_{m+1}}{\lambda_{m}}\right|\right)\frac{\left\lVert\Psi_{\leq m}^{\text{T}}V\Phi\right\rVert_{2}\cdot\left\lVert\Psi_{>m}^{\text{T}}V\Phi\right\rVert_{2}}{\left\lVert V\Phi\right\rVert_{2}},$ (33)
@@
$\Theta(\mathcal{R}(V^{\prime}),\mathcal{R}(\Psi_{\leq m}))<\Theta(\mathcal{R}(V),\mathcal{R}(\Psi_{\leq m}))$ (34)
@@
Proof: The left-hand side of the inequality (34) is the largest singular value of the matrix $T_{V^{\prime}}$ which we write as
$T_{V^{\prime}}$ $=\Psi_{>m}^{\text{T}}(AV\Phi+E)\left[\Psi_{\leq m}^{\text{T}}\left(AV\Phi+E\right)\right]^{-1}$ (35)
@@
$\sigma_{1}(T_{V^{\prime}})$ $\leq\sigma_{1}(\Lambda_{>m}V_{>m}\Phi+E_{>m})\cdot\sigma_{1}\left((\Lambda_{\leq m}V_{\leq m}\Phi+E_{\leq m})^{-1}\right)$ (36)
@@
$\sigma_{1}\left(\Lambda_{>m}V_{>m}\Phi+E_{>m}\right)$ $\leq\sigma_{1}\left(\Lambda_{>m}V_{>m}\Phi\right)+\left\lVert E_{>m}\right\rVert_{2}$ (37)
@@
$\sigma_{m}\left(\Lambda_{\leq m}V_{\leq m}\Phi+E_{\leq m}\right)$ $\geq\sigma_{m}\left(\Lambda_{\leq m}V_{\leq m}\Phi\right)-\left\lVert E_{\leq m}\right\rVert_{2}$ (38)
@@
$\sigma_{1}(T_{V^{\prime}})\leq\frac{|\lambda_{m+1}|\cdot\left\lVert V_{>m}\Phi\right\rVert_{2}+\left\lVert E_{>m}\right\rVert_{2}}{|\lambda_{m}|\cdot\sigma_{m}(V_{\leq m}\Phi)-\left\lVert E_{\leq m}\right\rVert_{2}}.$ (39)
@@
$\sigma_{1}(T_{V})$ $=\sigma_{1}\left((V_{>m}\Phi)(V_{\leq m}\Phi)^{-1}\right)$ (40)
@@
$\frac{|\lambda_{m+1}|\cdot\left\lVert V_{>m}\Phi\right\rVert_{2}+\left\lVert E_{>m}\right\rVert_{2}}{|\lambda_{m}|\cdot\sigma_{m}(V_{\leq m}\Phi)-\left\lVert E_{\leq m}\right\rVert_{2}}<\frac{\left\lVert V_{>m}\Phi\right\rVert_{2}}{\left\lVert V_{\leq m}\Phi\right\rVert_{2}}.$ (41)
@@
$\left\lVert V_{\leq m}\Phi\right\rVert_{2}\left(|\lambda_{m+1}|\left\lVert V_{>m}\Phi\right\rVert_{2}+\left\lVert E_{>m}\right\rVert_{2}\right)<\left\lVert V_{>m}\Phi\right\rVert_{2}\left(|\lambda_{m}|\sigma_{m}(V_{\leq m}\Phi)-\left\lVert E_{\leq m}\right\rVert_{2}\right),$ (42)
@@
$\left\lVert V_{\leq m}\Phi\right\rVert_{2}\left\lVert E_{>m}\right\rVert_{2}+\left\lVert V_{>m}\Phi\right\rVert_{2}\ \left\lVert E_{\leq m}\right\rVert_{2}$ (43)
@@
$\left\lVert V_{\leq m}\Phi\right\rVert_{2}\left\lVert E_{>m}\right\rVert_{2}+\left\lVert V_{>m}\Phi\right\rVert_{2}\ \left\lVert E_{\leq m}\right\rVert_{2}$ $\leq\left\lVert\begin{bmatrix}\left\lVert V_{\leq m}\Phi\right\rVert_{2}\\ \left\lVert V_{>m}\Phi\right\rVert_{2}\end{bmatrix}\right\rVert_{2}\left\lVert\begin{bmatrix}\left\lVert E_{>m}\right\rVert_{2}\\ \left\lVert E_{\leq m}\right\rVert_{2}\end{bmatrix}\right\rVert_{2}$ (44)
@@
$\left\lVert E\right\rVert_{2}<\left(\frac{|\lambda_{m}|}{\kappa_{2}(V_{\leq m}\Phi)}-|\lambda_{m+1}|\right)\frac{\left\lVert V_{\leq m}\Phi\right\rVert_{2}\left\lVert V_{>m}\Phi\right\rVert_{2}}{\left\lVert V\Phi\right\rVert_{2}}$ (45)
where $\kappa_{2}(V_{\leq m}\Phi)=\sigma_{1}(V_{\leq m}\Phi)/\sigma_{m}(V_{\leq m}\Phi)$. Finally dividing the preceding inequality by $|\lambda_{m}|$ we arrive at (33). $\square$
@@
If the eigenvectors of interest $\psi_{1},\ldots,\psi_{m}$ have rank less than or equal to $\bm{r}$, the truncation error $E^{\texttt{mv}}$ resulting from the matrix-vector products in line 3 of the truncated subspace iteration (algorithm 2) naturally decreases as the method converges. Similarly, the rank truncation in line 4 of the low-rank Rayleigh-Ritz procedure (algorithm 3) introduces error that also decreases as the subspace iteration converges. This is because $v_{j}$’s converge to $\psi_{j}$’s, for j=1,2,,mj=1,2,...,m, which are naturally orthogonal for symmetric $A$. As a result, the linear combinations in the Rayleigh-Ritz step involve fewer terms of significant magnitude, thereby reducing truncation error.
5 Computational cost and implementation
In this section, we estimate the computational cost and discuss several practical issues that are important for efficient implementations of subspace iteration with rank truncation.
@@ 5.2 Computing linear combination of low-rank TTs
In the Rayleigh-Ritz step (algorithm 3), approximating each linear combination in line 4 using the TT-SVD approach involves applying a truncation after each addition, resulting in m1m-1 truncations. Each truncation has a computational cost of $\mathcal{O}(dnr^{3})$ and the ordering of the summands can affect the final result. Alternatively, approximating the entire sum directly on the tangent space of the TT manifold at $z_{k}$ requires a single orthogonal projection, which costs $\mathcal{O}(dnr^{3})$. For a large number of summands $m$, the tangent space projection becomes significantly more efficient.
@@ 6 Numerical results
In this section, we present several numerical examples to demonstrate the efficiency and effectiveness of the rank truncated subspace iteration and compare it with both the rank truncated Lanczos algorithm and the DMRG algorithm for computing several algebraically smallest eigenvalues. To initialize the linear transformation defining the Chebyshev polynomial (12) we estimate the largest eigenvalue using a few rank-truncated Lanczos iterations with safeguards as outlined in [40] and set $b$ equal to such approximation. We initialize a=1a=-1 and update $a$ to be the largest Ritz value after each subspace iteration.
@@ 6.1 Test problems
    $H=\sum_{j=1}^{L}-J\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+\sigma_{j}^{z}\sigma_{j+1}^{z}\right)-h\sigma_{j}^{z},$ (46)
@@
    $\Delta=-\left(D\otimes I\otimes\cdots\otimes I+\cdots+I\otimes\cdots\otimes I\otimes D\right),$ (47)
@@ A Hamiltonian with a Hénon-Heiles potential can be written as
    $H=-\frac{1}{2}\Delta+\frac{1}{2}\sum_{k=1}^{d}q_{k}^{2}+\mu\sum_{k=1}^{d-1}\left(q_{k}^{2}q_{k+1}-\frac{1}{3}q_{k+1}^{3}\right),$ (48)
@@
Each of the test problem Hamiltonians are written above in the canonical polyadic (CP) format (2). In order to apply the algorithms described above, each Hamiltonian is first converted from CP format to TT-matrix format.
@@ 6.2 Comparing subspace iteration and Lanczos method
We compared the rank truncated subspace iteration (algorithm 1) and Lanczos method (algorithm 5) for computing the first several smallest eigenvalues of each test problem (46)-(48). For (46) we set $L=10$ resulting in $H$ with size $2^{10}\times 2^{10}$, small enough to compute eigenvalues and eigenvectors using standard methods, which serve as benchmarks for evaluating the low-rank methods. fig. 2 shows the error versus iteration for the first $5$ Ritz values computed by the low-rank subspace iteration and the low-rank Lanczos method. While the convergence Lanczos method stagnates around $10^{-3}$, the subspace iteration achieves convergence to machine precision. We also ran the Lanczos method without rank truncation to obtain the orthogonal Krylov basis vectors that the rank-truncated Lanczos method aims to approximate. In fig. 2(c), we plot the maximum of the TT-rank vector required to represent the first $20$ Krylov basis vectors with accuracy $10^{-10}$. After just two iterations, this rank exceeds $30$ indicating that these Krylov basis vectors can not be efficiently represented with small TT rank. Furthermore, Krylov methods are known to be sensitive to large errors in the early iterations, indicating that the stagnation in convergence arises from the inability to accurately represent the orthogonal Krylov basis vectors as low-rank TTs. In contrast, the subspace iteration does not encounter convergence issues due to rank truncation. Unlike the Krylov basis vectors, the ground state and first excited states can be effectively represented as low-rank TTs, as shown in the right panel of fig. 2. Moreover the subspace iteration is robust to large truncation errors in the initial iterations.
@@ 6.3 Comparing subspace iteration and DMRG
We now compare rank-truncated subspace iteration with DMRG on spin-1/2 and spin-1 Heisenberg Hamiltonian (46) with $L=100$, no external magnetic field ($h=0$), J=1J=-1 and periodic boundary conditions. We tested the DMRG implementation block2 [39] and the Matlab TT-Toolbox implementation [6]. Both converged to the same energies and the results reported below were obtained with block2. We computed the ground state and first excited state energies (two algebraically smallest eigenvalues) using subspace iteration and DMRG method and plot the energies versus CPU-time in fig. 3. In both cases we used a maximum rank of 100 for all vectors and both methods are initialized with the same random initial vectors. For the spin-1/2 system, the DMRG method and subspace iteration converge to the same eigenvalues with DMRG converging significantly faster. For the spin-1 system, DMRG also converges faster but gets stuck in a local minimum, a behavior that has been previously reported [37]. In contrast, subspace iteration achieves lower energies, with the ground state energy converging to the reference value. Note that the subspace correction for DMRG proposed in [37] converges to the ground state for the spin-1 system considered here. The correction enriches the local MPS basis of the current iterate $v$ with components of $Av$, similar to power iteration which uses $\mathfrak{T}_{\bm{r}}(Av)$ as the updated vector.